1) How many moles of carbon monoxide would be required to react with 3.24 moles of iron (III) oxide?                           Fe2O3(s)+ 3 CO(g) 🡪 2 Fe(g) + 3 CO2(g) 


1
Expert's answer
2021-02-24T03:41:56-0500

Solution:

The balanced chemical equation:

Fe2O3(s) + 3CO(g) → 2Fe(g) + 3CO2(g)

According to the equation above: n(Fe2O3) = n(CO)/3


n(CO) = 3 × n(Fe2O3) = 3 × 3.24 mol = 9.72 mol

n(CO) = 9.72 mol


Answer: 9.72 moles of carbon monoxide (CO) would be required.